# Investigating Experiential Effects in Online Chess using a Hierarchical Bayesian Analysis

### v1

@@ 5 Model Validation
The previous sections demonstrated that, under the proposed hierarchical regression model, there is little evidence for a consistent winner-/loser-effect across the population of online chess players studied. However, our model is very much a simplification of the complex nature of such data. One important consideration which are model does not account for is the temporal nature of this data. Many of the focal players in these datasets have several years of games in the dataset. It is possible that potential experiential effects can change over time, which we would fail to capture in this model. There is the potential for strong temporal dependence in this data which we do not account for in our simplified model. Below we consider several methods to evaluate the fit of our proposed model in the context of these potential issues. We demonstrate that while our model may somewhat suffer from ignoring the clear temporal structure in this data, the current model can capture much of the dependence in this data.
@@
[…]
where $\theta$ is the vector of parameters in the model. We can then draw replicated data from this distribution, namely we can simulate the result of a given match, given the covariates and each posterior draw of the model parameters. A failure of these replications to match the observed data can identify potential flaws in the model. These posterior predictive checks are often done using some generated quantity from the original dataset. Here, a natural quantity to consider is the number of games won by a focal player in the observed data. An extension of this is to consider the generated quantity corresponding to the rating evolution of a player over replicated datasets. This is the quantity we will use for posterior predictive checking below.
@@
We fit our previous proposed model with $n=1$ on the previous 1000 games to the final 1000 games for the GM cohort. As we use the standard number of chains (4) and sampling iterarions (1000) in CmdStanR, this gives us 4000 posterior draws for each parameter. We then use each set of draws to simulate the results of each game in the last 1000 games. This gives us 4000 replications of the 1000 games played by each focal player and the corresponding evolution of their Glicko-2 score. We show these replications for 2 GMs in Figure 7. While these draws show large uncertainty we see that they capture the overall trend well, in particular capturing the steady increase in rating shown in Figure 7(b). This indicates that our model is able to predict the future temporal evolution of the ability of players well.
@@
Fig. 7: Posterior predictive distribution of Glicko-2 ratings for 2 GMs. We show the true Lichess rating in red, with 4000 draws of the predicted rating evolution. The mean of these draws is shown in the solid black line.
Permuting the Game Results
As an alternative approach to assessing the robustness of our model, we consider permuting the data. A similar approach was considered by [11], to examine evidence for a hot hand effect in hockey. For computational considerations, we perform this analysis with a subset of the complete bullet data. We use 12 players from the 1700-1900 cohort and 10 players from the GM cohort, taking the final 1000 games from each. We randomly permute the results within each focal players data 1000 times, giving us 1000 permuted datasets. For each of these we fit our proposed model with $n=1$. We take the posterior mean for each of our parameters from each of these datasets, to construct a permutation distribution for each parameter. In Figure 8(a) we show the distribution of the posterior means under the permuted data, for the 1700-1900 cohort, for the estimated individual winner-/loser-effect for the 20 focal players. The intervals shown correspond to 66% and 95% of these permutation distributions. The estimated posterior means for the true data are shown in red. In each case these permuted datasets give similar estimated effects to those found with the real data, and all estimated effects agree with these null distributions.
We show the same analysis for the 10 players in the GM cohort in Figure 8(b). We see similar distributions for the individual effects under these permutations, although we see that the permutation intervals do not describe the true effects as well as in the 1700-1900 cohort. This is a general pattern we see where it appears that our model is less appropriate as the ability of players increases, and we see this as the rating increases. The estimated effects for the true data are similar to those estimated under the permutations. We note that when we repeat this analysis for the fixed effects $\gamma_{1}$ and $\gamma_{2}$ the distribution of these effects under the permutations is clearly different from the estimated effects, for all cohorts, as would be expected.
(a) Permutation estimates for 20 players in 1700-1900 Cohort.
(b) Permutation estimates for 10 players in the GM cohort.
Fig. 8: Posterior mean estimates for winner/loser-effects after permuting the game data. For each the true posterior mean estimate is shown in red, with the intervals representing 66% and 95% of the permutation distributions.
@@ 6 Summary, Future Directions
In this work we have proposed a Bayesian hierarchical logistic regression model to investigate the presence of experiential effects in online chess. having fit this model to cohorts of players of varying ability we find that there is no evidence for the presence of global winner/loser-effects, as commonly defined in the literature. However, a small proportion of individual players from all cohorts show evidence of possessing such effects. This is consistent across multiple formats of chess and the number of previous games that this effect could persist over. While we do not believe that our model perfectly describes the dynamics of chess, our model checking procedures indicate that our model appears reasonable and captures many of the key characteristics of this data. Our results indicate that a more flexible model may be required for players of higher ability, such as GMs. This would need to be investigated further, and could be related to the larger impact of the covariates (color and rating difference) in games between better players.
There are many potential future directions which could expand on this work. In shorter chess games it is common to attempt to ``flag" your opponent, aiming to beat them by forcing them to run out of time. A large proportion of bullet games are won this way (24% of bullet games in the GM cohort) and results of this form may have a different impact on future performance than other endings. We have also not attempted to separate out winner and loser effects or allow them to have different magnitudes in the current formulation. The current model also does not consider draws (which are rare in this data) and the potential impact they could have on future performance.
@@
Please insert acknowledgments of the assistance of colleagues or similar notes of appreciation here.
\@noitemerr
Please insert information concerning research grant support here (institution and grant number). Please provide for each funder the funder’s DOI according to https://doi.crossref.org/funderNames?mode=list.
@@
Here we include some additional model fitting results which were omitted from the main text. In particular, we show selected model estimates and model diagnostics for the other cohorts considered. We also investigate the effect of using varying numbers of games for the historic influence.
@@
Fig. A1: The estimated change in win probability for the other two cohorts we consider.
@@
Fig. A2: Posterior predictive distribution of Glicko-2 ratings for 3 players in different rating cohorts. We show the true Lichess rating in red, with 4000 draws of the predicted rating evolution. The mean of these draws is shown in the solid black line.
@@
Fig. A3: Posterior estimates for the key global parameters $\gamma_{1},\gamma_{2},\mu_{\beta}$ as we vary $n$ in our fitted model.
@@
Fig. A4: Estimated Player Effects for the 1700-1900 and GM cohorts.

### v2

@@ 5 Model Validation
The previous sections demonstrated that, under the proposed hierarchical regression model, there is little evidence for a consistent winner-/loser-effect across the population of online chess players studied. However, our model is very much a simplification of the complex nature of such data. One important consideration which our model does not account for is the temporal nature of this data. Many of the focal players in these datasets have several years of games in the dataset. It is possible that potential experiential effects can change over time, which we would fail to capture in this model. There is the potential for strong temporal dependence in this data which we do not account for in our simplified model. Below we consider one method to evaluate the fit of our proposed model in the context of these potential issues. We include an additional model checking procedure in the appendix, examining permutation distributions of the game results. We also include there a simulation study, indicating roughly how many games would be required to detect experiential effects of varying sizes. We demonstrate that while our model may somewhat suffer from ignoring the clear temporal structure in this data, the current model can capture much of the dependence in this data.
@@
$p(y^{\text{rep}}|y)=\int p(y^{\text{rep}}|\theta)p(\theta|y)d\theta,$ (4)
where $\theta$ is the vector of parameters in the model. We can then draw replicated data from this distribution, namely we can simulate the result of a given match, given the covariates and each posterior draw of the model parameters. A failure of these replications to match the observed data can identify potential flaws in the model. These posterior predictive checks are often done using some generated quantity from the original dataset. Here, a natural quantity to consider is the number of games won by a focal player in the observed data. An extension of this is to consider the generated quantity corresponding to the rating evolution of a player over replicated datasets, simulating each game result and then using these outcomes to construct the game history for future games. This is the quantity we will use for posterior predictive checking below.
@@
We fit our previous proposed model with $n=1$ on the previous 1000 games before then constructing a posterior predictive distribution of the Glicko-2 rating of each player using the final 1000 games for the GM cohort. As we use the standard number of chains (4) and sampling iterarions (1000) in CmdStanR, this gives us 4000 posterior draws for each parameter. We then use each set of draws to simulate the results of each game in the last 1000 games. This gives us 4000 replications of the 1000 games played by each focal player and the corresponding evolution of their Glicko-2 score. We show these replications for 2 GMs in Figure 7. While these replications show large uncertainty we see that they capture the overall trend well, in particular capturing the steady increase in rating shown in Figure 7(b). This indicates that our model is able to predict the future temporal evolution of the ability of players well. We include an additional model checking approach, examining the results of our model when we permute the game results in the appendix, along with this posterior predictive check for players from each of the other cohorts.
@@
Figure 7: Posterior predictive distribution of Glicko-2 ratings for 2 GMs. We show the true Lichess rating in red, with 4000 draws of the predicted rating evolution. The mean of these draws is shown in the solid black line.
@@ 6 Summary, Future Directions
In this work we have proposed a Bayesian hierarchical logistic regression model to investigate the presence of experiential effects in online chess. Having fit this model to cohorts of players of varying ability we find that there is no strong evidence for the presence of global winner/loser-effects, as commonly defined in the literature. However, a small proportion of individual players from all cohorts show evidence of possessing such effects. This is consistent across multiple formats of chess and the number of previous games that this effect could persist over. Our model is a general Bayesian model which can easily be modified to other data of this form, providing a tool for estimating both individual and population level experiential effects. While we do not believe that our model perfectly describes the dynamics of chess, our model checking procedures indicate that our model appears reasonable and captures many of the key characteristics of this data. Our results indicate that a more flexible model may be required for players of higher ability, such as GMs, where we saw some limited evidence for $\mu_{\beta}>0$. This would need to be investigated further, and could be related to the larger impact of the covariates (color and rating difference) in games between better players. Another concern which arises when examining these top players is that the opponent pool is significantly smaller. As such, the assumptions of playing random opponents may not be appropriate for particularly strong players. Modifying our approach to account for non-random opponents is an important future extension, along with examining the reliability of the history covariate we use here. A key advancement of the Glicko-2 rating system is the ability to incorporate individual rating deviation when examining the rating difference between players [25]. This gives a more reliable measure of the difference in ability between opponents, however this rating deviation is not available from the Lichess API. Being able to obtain and use this deviation could significantly improve our model.
There are many other potential future directions which could expand on this work. In shorter chess games it is common to attempt to ``flag" your opponent, aiming to beat them by forcing them to run out of time. A large proportion of bullet games are won this way (24% of bullet games in the GM cohort) and results of this form may have a different impact on future performance than other endings. We have also not attempted to separate out winner and loser effects or allow them to have different magnitudes in the current formulation. The current model also does not consider draws (which are rare in this data) and the potential impact they could have on future performance. Similarly, while we consider Bullet and Blitz games separately here we could also use both for predicting the outcome of either format, although it is very uncommon for players to switch between formats in a given session. More broadly, given the general nature of the model proposed here, we hope to also apply it to recent large animal interaction datasets, where the presence of experiential effects remains hotly debated.
@@
Adam Gee acknowledges the support of an NSERC CGRS M scholarship and Owen Ward acknowledges support from an NSERC Discovery Grant RGPIN-2023-05335.
@@
Here we include some additional model fitting results which were omitted from the main text. In particular, we show selected model estimates and model diagnostics for the other cohorts considered along with an additional model checking procedure and simulation study. We also investigate the effect of using varying numbers of games for the historic influence.
@@
Figure A1: The estimated change in win probability for the other two cohorts we consider.
@@
Figure A2: Posterior predictive distribution of Glicko-2 ratings for 3 players in different rating cohorts. We show the true Lichess rating in red, with 4000 draws of the predicted rating evolution. The mean of these draws is shown in the solid black line.
@@
Figure A3: Posterior estimates for the key global parameters $\gamma_{1},\gamma_{2},\mu_{\beta}$ as we vary $n$ in our fitted model.
Permuting the Game Results
As an alternative approach to assessing the robustness of our model, we consider permuting the data. A similar approach was considered by [11], to examine evidence for a hot hand effect in hockey. For computational considerations, we perform this analysis with a subset of the complete bullet data. We use 20 players from each of the 1700-1900 and GM cohorts, taking the final 1000 games from each. We randomly permute the results within each focal player's data 1000 times, giving us 1000 permuted datasets. For each of these we fit our proposed model with $n=1$. We then examine the distribution of the posterior means of the experiential effects that are generated from each of these model fits.
Figure A4 shows these ``permutation distributions'' for 20 players in the 1700-1900 cohort and 20 players in the GM cohort, with the central 66% and 95% intervals for each estimate of $\beta_{j}$ under these permutation distributions. For all players the true estimated experiential effects lie within these 95% permutation intervals. This aligns with our previous results, indicating that individual experiential effects are small and close to zero. We note that when we repeat this analysis for the fixed effects $\gamma_{1}$ and $\gamma_{2}$ the distribution of these effects under the permutations is clearly different from the estimated effects, for all cohorts, as would be expected.
(a) Permutation estimates for 20 players in 1700-1900 Cohort.
(b) Permutation estimates for 20 players in the GM cohort.
Figure A4: Posterior mean estimates for winner/loser-effects after permuting the game data. For each the true posterior mean estimate is shown in red, with the intervals representing 66% and 95% of the estimates obtained under the permutation distributions.
Simulation Study
Our previous analyses reveal that the experiential effects that we are interested in, if they exist, are likely small. In this case, even the large number of games analysed here may not be sufficient to detect such small effects. To investigate this further we construct a simulation study, examining when these effects can be identified from a cohort of 10 simulated focal players. To do this, we simulate game outcomes from the model proposed in Section 3. The values of the parameters in this model are given by:
  • $\gamma_{1}=0.091$
  • $\gamma_{2}=0.0038$
  • $\alpha_{j}\sim N(-0.075,0.025)$. This distribution approximates the estimates from Figure 6(a).
  • $\beta_{j}\sim N(\mu_{\beta},0.01)$. This variance is close to the posterior mean for $\tau_{2}$ estimated from the 1700-1900 cohort.
  • The rating difference for any given game is sampled from a Laplace distribution with mean 0 and scale 30. This closely resembles the actual rating differences in the 1700-1900 cohort.
  • For any given game, the simulated player had 0.5 probability of being white and 0.5 probability of being black.
We choose to only look at the previous game for history here ($n=1$). To investigate what effects we can detect we vary the true value of $\mu_{\beta}$ in this simulation along with the number of games played by each player, using $\mu_{\beta}=(-0.5,0.12,0.5)$ and $n_{\text{per player}}=(1,000,\;5,000,\;20,000)$ We chose two large global experiential effects in different directions and one small positive experiential effect, giving a range of possible true effects. Note that $\mu_{\beta}=0.12$ equates to an approximate average boost in win percentage by 3% when coming from a win for all players in the dataset. We chose a maximum of 20,000 games per player to be somewhat larger than the average number of games per player in the 1700-1900 cohort. We show the posterior estimates of $\mu_{\beta}$ and $\beta_{j}$ for each of these 9 simulation settings in Figure A5.
(a) The estimated values of $\mu_{\beta}$ for 3 different known effect sizes.
(b) The estimated experiential effects for 10 simulated players for known $\mu_{\beta}=0.12$.
Figure A5: 66% and 95% credible intervals for estimated effects for $\mu_{\beta}$ and $\beta_{j}$. In both figures the red vertical lines represent the true values of each respective parameter.
In each simulation study, the 95% credible intervals contain the true parameter values for both $\mu_{\beta}$ and each $\beta_{j}$. We note that when estimating individual $\beta_{j}$, we need a large number of games to get posterior intervals which do not contain 0 when the true value is non-zero. While a large number of games is required to detect these small effects, as shown in our model checking procedures, these small effects have little importance in describing the outcomes of these games.
@@
Figure A6: Estimated Player Effects for the 1700-1900 and GM cohorts.